Please help
2x + 6y + 3z = 24
5x − 3y − 5z = 3
4x + 3y + 2z = 33
What does x y and z =

Answer by stanbon(75887)

2x + 6y + 3z = 24
5x − 3y − 5z = 3
4x + 3y + 2z = 33
------
x = 6
y = -1
z = 6
===========
Cheers,
Stan H.
===========